OpenStudy (anonymous):

5 more:) 6x(x − 4) − 16x2 − (9x − 1)

OpenStudy (anonymous):

6x(x - 4) = 6x^2 - 24x, and - (9x - 1) = -9x + 1, So we have 6x2 - 24x - 16x2 -9x + 1 Combine like terms: -10x2 - 33x + 1

OpenStudy (llort):

DO YOUR OWN HOMEWORK!!!

OpenStudy (llort):

\[x = 1/20 (-33-\sqrt(1129))\] \[x = 1/20 (\sqrt(1129)-33)\]

OpenStudy (anonymous):

Thanks @Llort what is that^

OpenStudy (anonymous):

That's him giving you a bogus answer because he's angry that you're getting help on your homework in a forum that is designed to help people on their homework. Puzzling, really.

OpenStudy (anonymous):

Lol thank you hickninja i was thinkin the same thing

OpenStudy (llort):

those are the two points of x where y=0